OpenStudy (anonymous):

HELP ASAP MEDAL Which of the following functions is equivalent to the one shown below? y = -4x^2+56x-194 A. y-2 = -4(x-7)^2 B. y-194 = -4(x+7)^2 C. y= - (2x+4)(2x+48) D. y= -4x(x-14)

OpenStudy (anonymous):

@HatchetLover i'm going to ask for your help a lot lol

OpenStudy (anonymous):

y-2 = -4(x^2-14x+49) y-2 = -4x^2+56x-196 +2 +2 -------------------- y=-4x^2+56x-194

OpenStudy (anonymous):

so the answer is A

OpenStudy (anonymous):

gracias:)

OpenStudy (anonymous):

de nada(: